How much work does a supermarket checkout attendant do on a can of soup he pushes 0.610 m horizontally with a force of 4.80 N? Express your answer in joules and kilocalories. kcal
Answer #1

Solution)

We know, Work done,

W = F * d = 4.8* 0.610 = 2.928 =3Joules (Ans)

Now, we know

1 cal = 4.186 J

So,

W = 3 / 4.186 = 0.7166 calories=0.0007166 kcal (Ans)

=====

Good luck!:)